# Intuitionistic NF set theory

### v5

@@ 3. Finite sets
That formula is stratified, giving $X$ and $Y$ index 1, since the similarity relation can be defined in INF. Therefore induction is legal.
Base case, $X=\Lambda$. The only subset of $\Lambda$ is $\Lambda$, so any subset of $X$ is equal to $X$. That completes the base case.
@@
(21) $Y-\{c\}=A.$
Then $c\not\in A$. But $X=A\cup\{b\}$, and $c=f(b)\in Y\subseteq A$, so $c\in A$. That contradiction completes Case 3, and that completes the proof of the induction step. That in turn completes the proof of the theorem.
@@
$\land\ \forall u,v\in x\,(u\neq v\ {\rightarrow}\ u\cap v=\Lambda)$
@@
Proof. By induction on the finite set $x$. Base case, $x=\Lambda$. Then $\bigcup x=\Lambda$, which is finite.
@@
(22) $\bigcup x$ $=$ $\left(\bigcup y\right)\cup c$
Now $c$ is finite, since every member of $x$ is finite and $c\in x$. We have $\bigcup y\cap c=\Lambda$, since if $p$ belongs to both $\bigcup y$ and $c$, then for some $w\in y$ we have $p\in w\cap c$, contradicting the hypothesis that any two distinct members of $x=y\cup\{c\}$ are disjoint. Then $\bigcup y\cup c$ is finite, by Lemma 3.11. Then $\bigcup x$ is finite, by (22). That completes the induction step, and that completes the proof of the lemma.
@@
$SSC(x)\subseteq SSC(x\cup\{c\}).$
Proof. About 30 straightforward steps, which we choose to omit here.
@@
Proof. Let $X$ and $Y$ be two separable subsets of $A$. Let $u\in A$. By definition of separability, we have
@@
(23) $u\in X\cap Y\ \lor\ u\not\in X\cap Y.$
This can be proved by cases; there are four cases according to whether $u$ is in $X$ or not, and whether $u$ is in $Y$ or not. In each case, (23) is immediate. Hence $X\cap Y$ is a separable subset of $A$, as claimed. Similarly, $X\cup Y$ is a separable subset of $A$. That completes the proof of the lemma.
@@
Proof. The formula of the lemma is stratified, giving $x$ index 0, $B$ index 1, and $P$ index 1. Therefore we may proceed by induction on finite sets $B$. (Notice that the statement being proved by induction is universally quantified over $P$–that is important because in the induction step we need to substitute a different set for $P$; the proof does not work with $P$ a parameter.)
Base case, $B=\Lambda$. The conclusion $\forall x\in\Lambda\,x\in P$ holds since $x\in\Lambda$ is false.
@@
(25) $\neg\neg\,\forall x\in B\,(x\in P\ \land\ c\in P)$   by the definition of $Q$
@@
(26) $\neg\neg\,((\forall x\in B\,(x\in P))\ \land\ c\in P),$
which seems plausible as $x$ does not occur in ‘$c\in P$’. In fact we have the equivalence of (25) and 26), since $B$ is inhabited. (That was why we had to break the proof into cases according as $B$ is empty or inhabited.) Then indeed (26) follows. By the definitions of union and unit class we have
@@
That completes the induction step. That completes the proof of the lemma.
@@
Remark. We already know that separable subsets of a finite set are finite, and finite subsets of finite set are separable, but one cannot hope to prove every subset of a finite set is finite, because of sets like $\{x\in\{\Lambda\}:P\}$. That set is finite if and only if $P\ \lor\ \neg P$, by Lemma 3.4.
Proof. Let $X$ be a finite set, and $A\subseteq X$. By Lemma 3.19, if $A$ is a separable subset of $X$ then $A$ is finite. Double-negating that implication, if $A$ is not-not separable, then it is not-not finite. Hence, it suffices to prove that not-not $A$ is a separable subset of $X$. More formally, we must prove
(27) $\neg\neg\,X=A\cup(X-A)$
@@
That is (27), so that completes the proof of the lemma.
@@
Remark. Lemma 3.11 shows the double negation can be dropped if $x$ and $y$ are assumed to be disjoint. It cannot be dropped in general, as $\{a\}\cup\{b\}\in{\sf FINITE}$ implies $a=b\ \lor a\neq b$, so if we could drop the double negation in this lemma, then every set would have decidable equality.
Proof. The formula is stratified, so we can prove it by induction on finite sets $y$, for a fixed finite set $x$.
Base case, $y=\Lambda$. We have $x\cup\Lambda=x$, which is finite by hypothesis. That completes the base case.
@@
That completes the induction step. That completes the proof of the lemma.
@@
That completes the proof of the lemma.
4. Frege cardinals
The formula in the following definition is stratifiable, so the definition can be given in INF. Specifically, we can give $a$ index 0, $x$ and $z$ index 1, and $\kappa$ index 2. Then $\kappa^{+}$ gets index 2, so the successor function $\kappa\mapsto\kappa^{+}$ is a function in INF.
@@ 4. Frege cardinals
The set ${\mathbb{F}}$ of finite Frege cardinals is the least set containing ${\sf zero}=\{\Lambda\}$ and containing $\kappa^{+}$ whenever it contains $\kappa$ and $\kappa^{+}$ is inhabited. More precisely,
@@
Remarks. The formula defining ${\mathbb{F}}$ is stratified, so the definition can be given in INF. According to that definition, if there were a largest finite cardinal $\kappa$, then $\kappa^{+}$ would be the empty set, not Frege zero, which is $\{\Lambda\}$. So in that case, the successor of the largest finite cardinal $\kappa$ would not belong to ${\mathbb{F}}$, which does not contain $\Lambda$. Instead, in that case the result would be that successor does not map ${\mathbb{F}}\to{\mathbb{F}}$. Of course $\Lambda^{+}=\Lambda$, so once that happened, more applications of successor would do nothing more. Note also that in general a finite cardinal is not a finite set; rather, the members of a finite cardinal are finite sets.
@@
The formula in the definition is stratifiable, so the definition is legal. We will show that $Z$ is closed under the conditions defining ${\mathbb{F}}$. First, Frege zero =$=\{\Lambda\}$ is in $Z$, since $\Lambda$ is finite. To verify the second condition, assume $\kappa\in Z$ and $\kappa^{+}$ is inhabited; we must show $\kappa^{+}\in Z$. Let $u\in\kappa^{+}$. Then there exists $x\in\kappa$ and there exists $a$ such that $u=x\cup\{a\}$. Since $\kappa\in Z$, $x$ is finite. Then by definition of FINITE, $u$ is finite. That completes the proof that $Z$ satisfies the second condition. Hence ${\mathbb{F}}\subseteq Z$. That completes the proof of the lemma.
@@
$(\phi({\sf zero})\ \land\ \forall x\,(\phi(x)\ \land\exists u\,(u\in\ x^{+})\ {\rightarrow}\ \phi(x^{+}))\ {\rightarrow}\ \forall x\,\phi(x).$
Proof. $Z:=\{x:\phi(x)\}$ is definable and satisfies the closure conditions that define ${\mathbb{F}}$. Therefore ${\mathbb{F}}\subseteq Z$. That completes the proof of the lemma.
Remark. Note that when carrying out a proof by induction, during the induction step we get to assume that $x^{+}$ is inhabited.
@@
Remark. It would not do to use $\exists u$ instead of $\forall u$, since then $\Lambda$ would not be a cardinal, but we need to allow for that possibility. We note that Rosser’s definition is equivalent to ours.
@@
Proof. Lemma 4.5 justifies us in proving $\exists u\,(u\in\kappa)$ by induction on $\kappa$.
Base case: ${\sf zero}=\{\Lambda\}$ is inhabited.
Induction step: Suppose $\kappa^{+}$ is inhabited. Then $\kappa^{+}$ is inhabited. (We do not even need to use the induction hypothesis.) That completes the proof of the lemma.
@@
That formula can be stratified, since we have already shown that $x\sim y$ is definable in INF. Therefore the definition of $Z$ is legal.
We will show $Z$ contains Frege zero and is closed under Frege successor. $Z$ contains Frege zero since the only member of Frege zero is the empty set, and the only set in one-to-one correspondence with the empty set is $\Lambda$ itself.
@@
I say that $v\cup\{b\}=y$. Let $p\in y$. Then $p=f(q)$ for some $q\in x$, since $f$ maps $x$ onto $y$. By Lemma 4.4, since $x\in\kappa^{+}$, $x$ is finite. Since $x$ is finite, it has decidable equality by Lemma 3.3. Therefore $q=a\ \lor\ q\neq a$. If $q=a$ then $p=f(a)=b\in\{b\}$. If $q\neq a$ then since $q\in x=u\cup\{a\}$ and $q\in x$, we have $q\in u$. Then by definition of $v$, $p=f(q)\in v$. Therefore $p\in v\cup\{b\}$. Since $p$ was an arbitrary member of $y$, we have proved $y\subseteq v\cup\{b\}$. But $v\cup\{b\}\subseteq y$ is immediate, since $v\subseteq y$ and $b\in y$. Therefore $v\cup\{b\}=y$, as claimed.
Since $v\in\kappa$, it follows that $y\in\kappa^{+}$ as desired. Thus $Z$ is closed under Frege successor. By the definition of ${\mathbb{F}}$, we have ${\mathbb{F}}\subseteq Z$. That completes the proof of the lemma.
@@
Proof. By induction on $\kappa$. Similarity is defined by a stratified formula, so induction is legal. The base case is immediate as Frege zero has only one member. For the induction step, let $x$ and $y$ belong to $\kappa^{+}$. Then there exist $u,v,a,b$ such that $u,v\in\kappa$ and $a\not\in u$ and $b\not\in v$ and $x=u\cup\{a\}$ and $y=v\cup\{b\}$. By the induction hypothesis, there is a one-to-one correspondence $g:u\to v$. We define $f:x\to y$ by
@@
Therefore $f$ is one-to-one, as claimed. Therefore $x\sim y$. That completes the induction step. That completes the proof of the lemma.
@@
$Nc(x)=\{u:u\sim x\}.$
Then the inhabited cardinals, that is, the inhabited members of $NC$, are exactly the sets of the form $Nc(x)$ for some $x$.
@@
For all $x$, $x\in Nc(x)$
lemma By Lemma 2.11, we have $x\sim x$. Then $x\in Nc(x)$ by Definition 4.10. That completes the proof of the lemma.
@@
$Nc(x)=Nc(y)$ if and only if $x\sim y$.
Proof. By Lemma 2.11, which says that the relation $\sim$ is an equivalence relation.
@@
$c\not\in x\ {\rightarrow}\ Nc(x\cup\{c\})=Nc(x)^{+}$.
Proof. By extensionality, it suffices to show that the two sides have the same members. That is, we must show, under the assumption $c\not\in x$,
@@
Ad right-to-left: Suppose $b\not\in v$ and $v\sim x$ and $u=v\cup\{b\}$. Let $f:v\to x$ be a similarity, and extend it to $g$ defined by $g=f\cup\{\langle b,c\rangle\}$. Then $g$ is a similarity from $v\cup\{b\}$ to $x\cup\{c\}$. That completes the right-to-left direction. f Ad left-to-right: Suppose $f:u\to x\cup\{c\}$ is a similarity. Since $f$ is onto, there exists $b\in x$ with $f(b)=c$. Let $v=u-\{b\}$. Use this $b$ and $v$ on the right. Then $g=f-\{\langle b,c\rangle\}$ is a similarity from $v$ to $x$. It remains to show $u=v\cup\{b\}=(u-\{b\})\cup\{b\}$. That is,
$z\in u\ {\rightarrow}\ z\in u\ {\rightarrow}\ z\neq b\ \lor\ z=b.$
Let $z\in u$. Since $f$ is a similarity from $u$ to $x\cup\{c\}$, there is a unique $y\in x\cup\{c\}$ such that $\langle z,y\rangle\in f$. Then $z=b{\ \leftrightarrow\ }y=c$. Since $c\not\in x$, and $y\in x\cup\{c\}$, $y=c\ \lor\ y\neq c$. Therefore $z=b\ \lor z\neq b$, as desired. Note that it is not necessary that $z$ have decidable equality. That completes the left-to-right direction. That completes the proof of the lemma.
@@
$Nc(\Lambda)={\sf zero}$.
Proof. By Definition 4.2, ${\sf zero}=\{\Lambda\}$. By definition, $Nc(\Lambda)$ contains exactly the sets similar to $\Lambda$. By Lemma 2.12, $\Lambda$ is the only set similar to $\Lambda$. Therefore $Nc(\Lambda)=\{\Lambda\}$. Then $Nc(\Lambda)={\sf zero}$ since both are equal to $\{\Lambda\}$. That completes the proof of the lemma.
@@
Remark. Note that $\kappa$ is not assumed to be a finite cardinal, or even a cardinal. Successor cannot take the value ${\sf zero}=\{\Lambda\}$ on any set. We will need that generality at a certain point.
Proof. By definition the members of $\kappa^{+}$ are exactly the sets of the form $x\cup\{a\}$ with $x\in\kappa$ and $a\not\in x$. (That is true whether or not there are any such members.) But if $\kappa^{+}$ is inhabited, then there is at least one such member, and each such member $x\cup\{a\}$ is inhabited, since it contains $a$. That completes the proof of the lemma.
@@
Remark. This does not depend on the finiteness or not-finiteness of ${\mathbb{F}}$. If ${\mathbb{F}}$ is finite then eventually $\kappa^{+}$ is $\Lambda$, rather than ${\sf zero}$, which is $\{\Lambda\}$, so even in that case Frege zero does not occur as a successor.
Proof. If $\kappa^{+}=\{\Lambda\}$ then $\kappa^{+}$ is inhabited, but contains no inhabited set, contradicting Lemma 4.15. That completes the proof of the lemma.
@@
Proof. The set $Z=\{\kappa\in{\mathbb{F}}:\kappa={\sf zero}\ \lor\ \exists\mu\,(\mu\in{\mathbb{F}}\ \land\ \kappa=\mu^{+})\}$ is definable in INF, since its defining formula is stratified. $Z$ contains Frege zero and is closed under successor. Therefore, by definition of ${\mathbb{F}}$, ${\mathbb{F}}\subseteq Z$. That completes the proof of the lemma.
@@
Proof. Let $W$ be one of the sets whose intersection defines ${\mathbb{F}}$, i.e., $W$ contains ${\sf zero}$ and is closed under inhabited successor. Then $W$ contains ${\sf zero}$. Since $W$ was arbitrary, ${\sf zero}\in{\mathbb{F}}$. That completes the proof.
@@
Proof. Suppose $\kappa\in{\mathbb{F}}$ and $\kappa^{+}$ is inhabited. Let $W$ be one of the sets whose intersection defines ${\mathbb{F}}$, i.e., $W$ contains ${\sf zero}$ and is closed under inhabited successor. By induction on $\kappa$, we can prove $\kappa\in W$. Since $W$ is closed under inhabited successor, and $\kappa\in W$, and $\kappa^{+}$ is inhabited, we have $\kappa^{+}\in W$. Since ${\mathbb{F}}$ is the intersection of all such sets $W$, and $\kappa^{+}$ belongs to every such $W$, we have $\kappa^{+}\in{\mathbb{F}}$ as desired. That completes the proof of the lemma.
@@
$\Lambda\in{\sf zero}$   by the definition of ${\sf zero}$
${\sf zero}\not\in{\sf zero}$   since ${\sf zero}=\{\Lambda\}$ and ${\sf zero}\neq\Lambda$
$\Lambda\cup\{{\sf zero}\}\in{\sf zero}^{+}$   by definition of successor
@@
$\forall x\in{\sf FINITE}\,(Nc(x)\in{\mathbb{F}}).$
Proof. The formula to be proved is stratified, so we can prove it by induction on finite sets.
Base case: By Lemma 4.14, $Nc(\Lambda)={\sf zero}$. By Lemma 4.19, ${\sf zero}\in{\mathbb{F}}$.
Induction step: Let $x\in{\sf FINITE}$ and $c\not\in x$. Consider $Nc(x\cup\{c\})$, which by Lemma 4.13 is $Nc(x)^{+}$. By the induction hypothesis, $Nc(x)\in{\mathbb{F}}$. By definition of ${\mathbb{F}}$, $Nc(x)^{+}\in{\mathbb{F}}$. That completes the induction step. That completes the proof of the lemma.
@@
Proof. By induction we prove
$\forall m\,(m\in{\mathbb{F}}\ {\rightarrow}\ \exists u\,(u\in m).$
The formula is stratified, giving $u$ index 0 and $m$ index 1. For the base case, ${\sf zero}=\{\Lambda\}$ by definition, so ${\sf zero}$ is inhabited. For the induction step, we always suppose $m^{+}$ is inhabited, so there is nothing more to prove. That completes the proof of the lemma. To put the proof directly: the set of inhabited members of ${\mathbb{F}}$ contains ${\sf zero}$ and is closed under inhabited successor, so it contains ${\mathbb{F}}$.
@@
Proof. Let $a$ be finite and $a\sim b$. Let $\kappa=Nc(a)$. Then

### v6

@@ 3 Finite sets
That formula is stratified, giving $X$ and $Y$ index 1, since the similarity relation can be defined in $\mathit{i}\mkern 0.4mu\mathrm{NF}$. Therefore induction is legal.
Base case, $X=\emptyset$. The only subset of $\emptyset$ is $\emptyset$, so any subset of $X$ is equal to $X$. That completes the base case.
@@
$Y-\{c\}=A.$ (21)
Then $c\not\in A$. But $X=A\cup\{b\}$, and $c=f(b)\in Y\subseteq A$, so $c\in A$. That contradiction completes Case 3, and that completes the proof of the induction step. ∎
@@
$\land\ \forall u,v\in x\,(u\neq v\ {\rightarrow}\ u\cap v=\emptyset)$
@@
Proof.
By induction on the finite set $x$.
Base case, $x=\emptyset$. Then $\bigcup x=\emptyset$, which is finite.
@@
$\bigcup x$ $=$ $\left(\bigcup y\right)\cup c$ (22)
Now $c$ is finite, since every member of $x$ is finite and $c\in x$. We have $\bigcup y\cap c=\emptyset$, since if $p$ belongs to both $\bigcup y$ and $c$, then for some $w\in y$ we have $p\in w\cap c$, contradicting the hypothesis that any two distinct members of $x=y\cup\{c\}$ are disjoint. Then $\bigcup y\cup c$ is finite, by Lemma 3.11. Then $\bigcup x$ is finite, by (22). That completes the induction step. ∎
@@
$\mathcal{P}_{s}(x)\subseteq\mathcal{P}_{s}(x\cup\{c\}).$
Proof.
About 30 straightforward steps, which we choose to omit here. ∎
@@
Proof.
Let $X$ and $Y$ be two separable subsets of $A$. Let $u\in A$. By definition of separability, we have
@@
$u\in X\cap Y\ \lor\ u\not\in X\cap Y.$ (23)
This can be proved by cases; there are four cases according to whether $u$ is in $X$ or not, and whether $u$ is in $Y$ or not. In each case, (23) is immediate. Hence $X\cap Y$ is a separable subset of $A$, as claimed. Similarly, $X\cup Y$ is a separable subset of $A$. ∎
@@
Proof.
The formula of the lemma is stratified, giving $x$ index 0, $B$ index 1, and $P$ index 1. Therefore we may proceed by induction on finite sets $B$. (Notice that the statement being proved by induction is universally quantified over PP–that is important because in the induction step we need to substitute a different set for $P$; the proof does not work with $P$ a parameter.)
Base case, $B=\emptyset$. The conclusion $\forall x\in\emptyset\,x\in P$ holds since $x\in\emptyset$ is false.
@@
$\neg\neg\,\forall x\in B\,(x\in P\ \land\ c\in P)$   by the definition of $Q$ (25)
@@
$\neg\neg\,((\forall x\in B\,(x\in P))\ \land\ c\in P),$ (26)
which seems plausible as $x$ does not occur in $c\in P$’. In fact we have the equivalence of (25) and 26), since $B$ is inhabited. (That was why we had to break the proof into cases according as $B$ is empty or inhabited.) Then indeed (26) follows. By the definitions of union and unit class we have
@@
That completes the induction step. ∎
@@
Remark. We already know that separable subsets of a finite set are finite, and finite subsets of finite set are separable, but one cannot hope to prove every subset of a finite set is finite, because of sets like $\{x\in\{\emptyset\}:P\}$. That set is finite if and only if P¬PP\ $\lor\ \neg$ P, by Lemma 3.4.
Proof.
Let $X$ be a finite set, and $A\subseteq X$. By Lemma 3.19, if $A$ is a separable subset of $X$ then $A$ is finite. Double-negating that implication, if $A$ is not-not separable, then it is not-not finite. Hence, it suffices to prove that not-not $A$ is a separable subset of $X$. More formally, we must prove
$\neg\neg\,X=A\cup(X-A)$ (27)
@@
That is (27). ∎
@@
[…]
Proof.
The formula is stratified, so we can prove it by induction on finite sets $y$, for a fixed finite set $x$.
Base case, $y=\emptyset$. We have $x\cup\emptyset=x$, which is finite by hypothesis. That completes the base case.
@@
That completes the induction step. ∎
@@
Before leaving this section, we shall state a technical lemma about similarities, arising from the details of the definitions of “maps” and “similarity”. The issue is that f:XYf:X\ $\rightarrow$}\ Y does not require that the domain of $f$ be exactly $X$; it is allowed to be larger. That is generally a good thing, as once we have defined $X$ and proved it maps $X$ to $Y$, it automatically maps subsets of $X$ to $Y$. But to be a similarity from $X$ to $Y$, the domain of $f$ must be exactly $X$ and the range exactly $Y$. The following lemma is the price we must pay for allowing the domain of $f$ to be larger in “maps”. Stating it here allows us to cite it, without revisiting this issue in future work.
Lemma 3.32.
Suppose f:ABf:A\ $\rightarrow$}\ B and $f$ is one to one and onto $B$. Let $R$ be the range of $f$. Suppose $R\subseteq B$ and the domain of $f$ is $A$. Then $f$ is a similarity from $A$ to $R$.
Proof.
We omit the proof, which takes 111 lines of Lean, because it is just a straightforward unwinding of the definitions involved. ∎
4 Frege cardinals
The formula in the following definition is stratifiable, so the definition can be given in $\mathit{i}\mkern 0.4mu\mathrm{NF}$. Specifically, we can give $a$ index 0, $x$ and $z$ index 1, and $\kappa$ index 2. Then $\kappa^{+}$ gets index 2, so the successor function $\kappa\mapsto\kappa^{+}$ is a function in $\mathit{i}\mkern 0.4mu\mathrm{NF}$.
@@ 4 Frege cardinals
The set $\mathbb{F}$} of finite Frege cardinals is the least set containing $\sf zero$$\{\emptyset\}$ and containing $\kappa^{+}$ whenever it contains $\kappa$ and $\kappa^{+}$ is inhabited. More precisely,
@@
Remarks. The formula defining $\mathbb{F}$} is stratified, so the definition can be given in $\mathit{i}\mkern 0.4mu\mathrm{NF}$. According to that definition, if there were a largest finite cardinal $\kappa$, then $\kappa^{+}$ would be the empty set, not Frege zero, which is $\{\emptyset\}$. So in that case, the successor of the largest finite cardinal $\kappa$ would not belong to $\mathbb{F}$}, which does not contain $\emptyset$. Instead, in that case the result would be that successor does not map ${\mathbb{F}}\to{\mathbb{F}}$. Of course $\emptyset^{+}=\emptyset$, so once that happened, more applications of successor would do nothing more. Note also that in general a finite cardinal is not a finite set; rather, the members of a finite cardinal are finite sets.
@@
The formula in the definition is stratifiable, so the definition is legal. We will show that $Z$ is closed under the conditions defining $\mathbb{F}$}. First, Frege zero $\{\emptyset\}$ is in $Z$, since $\emptyset$ is finite. To verify the second condition, assume $\kappa\in Z$ and $\kappa^{+}$ is inhabited; we must show $\kappa^{+}\in Z$. Let $u\in\kappa^{+}$. Then there exists $x\in\kappa$ and there exists $a$ such that $u=x\cup\{a\}$. Since $\kappa\in Z$, $x$ is finite. Then by definition of FINITE, $u$ is finite. That completes the proof that $Z$ satisfies the second condition. Hence ${\mathbb{F}}\subseteq Z$.
@@
$(\phi({\sf zero})\ \land\ \forall x\,(\phi(x)\ \land\exists u\,(u\in\ x^{+})\ {\rightarrow}\ \phi(x^{+})))\ {\rightarrow}\ \forall x\,\phi(x).$
Proof.
$Z:=\{x:\phi(x)\}$ is definable and satisfies the closure conditions that define $\mathbb{F}$}. Therefore ${\mathbb{F}}\subseteq Z$. ∎
Remark. When carrying out a proof by induction, during the induction step we get to assume that $x^{+}$ is inhabited.
@@
Remark. It would not do to use $\exists u$ instead of $\forall u$, since then $\emptyset$ would not be a cardinal, but allow for that possibility. We note that Rosser’s definition requires cardinals to be inhabited. In the work presented here, it makes no difference, as we work only with finite cardinals.
@@
Proof.
Lemma 4.5 justifies us in proving $\exists u\,(u\in\kappa)$ by induction on $\kappa$.
Base case: $\sf zero$$\{\emptyset\}$ is inhabited.
Induction step: Suppose $\kappa^{+}$ is inhabited. Then $\kappa^{+}$ is inhabited. (We do not even need to use the induction hypothesis.) ∎
@@
That formula can be stratified, since we have already shown that $x\sim y$ is definable in $\mathit{i}\mkern 0.4mu\mathrm{NF}$. Therefore the definition of $Z$ is legal.
We will show $Z$ contains Frege zero and is closed under Frege successor. $Z$ contains Frege zero since the only member of Frege zero is the empty set, and the only set in one-to-one correspondence with the empty set is $\emptyset$ itself.
@@
I say that $v\cup\{b\}=y$. Let $p\in y$. Then $p=f(q)$ for some $q\in x$, since $f$ maps $x$ onto $y$. By Lemma 4.4, since $\in\kappa^{+}, xx$ is finite. Since $x$ is finite, it has decidable equality by Lemma 3.3. Therefore $q=a\ \lor\ q\neq a$. If $q=a$ then $p=f(a)=b\in\{b\}$. If $q\neq a$ then since $q\in x=u\cup\{a\}$ and $q\in x$, we have $q\in u$. Then by definition of $v$, $p=f(q)\in v$. Therefore $p\in v\cup\{b\}$. Since $p$ was an arbitrary member of $y$, we have proved $y\subseteq v\cup\{b\}$. But $v\cup\{b\}\subseteq y$ is immediate, since $v\subseteq y$ and $b\in y$. Therefore $v\cup\{b\}=y$, as claimed.
Since $v\in\kappa$, it follows that $y\in\kappa^{+}$ as desired. Thus $Z$ is closed under Frege successor. By the definition of $\mathbb{F}$}, we have ${\mathbb{F}}\subseteq Z$. ∎
@@
Proof.
By induction on $\kappa$. Similarity is defined by a stratified formula, so induction is legal. The base case is immediate as Frege zero has only one member. For the induction step, let $x$ and $y$ belong to $\kappa^{+}$. Then there exist $u,v,a,b$ such that $u,v\in\kappa$ and $a\not\in u$ and $b\not\in v$ and $x=u\cup\{a\}$ and $y=v\cup\{b\}$. By the induction hypothesis, there is a one-to-one correspondence $g:u\to v$. We define $f:x\to y$ by
@@
Therefore $f$ is one-to-one, as claimed. Therefore $x\sim y$. That completes the induction step. ∎
@@
$|x|=\{u:u\sim x\}.$
Then the inhabited cardinals, that is, the inhabited members of $NC$, are exactly the sets of the form $|x|$ for some $x$.
@@
For all $x$, $x\in|x|$
lemma By Lemma 2.11, we have $x\sim x$. Then $x\in|x|$ by Definition 4.10. ∎
@@
$|x|=|y|$ if and only if $x\sim y$.
Proof.
By Lemma 2.11, which says that the relation $\sim$ is an equivalence relation. ∎
@@
$c\not\in x\ {\rightarrow}\ |x\cup\{c\}|=|x|^{+}$.
Proof.
By extensionality, it suffices to show that the two sides have the same members. That is, we must show, under the assumption $c\not\in x$,
$u\sim x\cup\{c\}$ ${\ \leftrightarrow\ }$ $\exists b,v\,(b\not\in v\ \land\ v\sim x\ \land\ u=v\cup\{b\}).$ (30)
Ad right-to-left: Suppose $b\not\in v$ and $v\sim x$ and $u=v\cup\{b\}$. Let $f:v\to x$ be a similarity, and extend it to $g$ defined by $g=f\cup\{\langle b,c\rangle\}$. Then $g$ is a similarity from $v\cup\{b\}$ to $x\cup\{c\}$. That completes the right-to-left direction.
f Ad left-to-right: Suppose $f:u\to x\cup\{c\}$ is a similarity. Since $f$ is onto, there exists $b\in x$ with $f(b)=c$. Let $v=u-\{b\}$. Use this $b$ and $v$ on the right. Then $g=f-\{\langle b,c\rangle\}$ is a similarity from $v$ to $x$. It remains to show $u=v\cup\{b\}=(u-\{b\})\cup\{b\}$. That is,
$z\in u\ {\rightarrow}\ z\in u\ {\rightarrow}\ z\neq b\ \lor\ z=b.$
@@
$|\emptyset|={\sf zero}$.
Proof.
By Definition 4.2, $\sf zero$$\{\emptyset\}$. By definition, $|\emptyset|$ contains exactly the sets similar to $\emptyset$. By Lemma 2.12, $\emptyset$ is the only set similar to $\emptyset$. Therefore $|\emptyset|=\{\emptyset\}$. Then $|\emptyset|={\sf zero}$ since both are equal to $\{\emptyset\}$. ∎
@@
Remark. Note that $\kappa$ is not assumed to be a finite cardinal, or even a cardinal. Successor cannot take the value $\sf zero$$\{\emptyset\}$ on any set.
Proof.
By definition the members of $\kappa^{+}$ are exactly the sets of the form $x\cup\{a\}$ with $x\in\kappa$ and $a\not\in x$. (That is true whether or not there are any such members.) But if $\kappa^{+}$ is inhabited, then there is at least one such member, and each such member $x\cup\{a\}$ is inhabited, since it contains $a$. ∎
@@
Remark. This does not depend on the finiteness or not-finiteness of $\mathbb{F}$}. If $\mathbb{F}$} is finite then eventually $\kappa^{+}$ is $\emptyset$, rather than $\sf zero$}, which is $\{\emptyset\}$, so even in that case Frege zero does not occur as a successor.
Proof.
If $\kappa^{+}=\{\emptyset\}$ then $\kappa^{+}$ is inhabited, but contains no inhabited set, contradicting Lemma 4.15. ∎
@@
Proof.
The set $Z=\{\kappa\in{\mathbb{F}}:\kappa={\sf zero}\ \lor\ \exists\mu\,(\mu\in{\mathbb{F}}\ \land\ \kappa=\mu^{+})\}$ is definable in $\mathit{i}\mkern 0.4mu\mathrm{NF}$, since its defining formula is stratified. $Z$ contains Frege zero and is closed under successor. Therefore, by definition of $\mathbb{F}$}, ${\mathbb{F}}\subseteq Z$. ∎
@@
Proof.
Let $W$ be one of the sets whose intersection defines $\mathbb{F}$}, i.e., $W$ contains $\sf zero$} and is closed under inhabited successor. Then $W$ contains $\sf zero$}. Since $W$ was arbitrary, ${\sf zero}\in{\mathbb{F}}$. ∎
@@
Proof.
Suppose $\kappa\in{\mathbb{F}}$ and $\kappa^{+}$ is inhabited. Let $W$ be one of the sets whose intersection defines $\mathbb{F}$}, i.e., $W$ contains $\sf zero$} and is closed under inhabited successor. By induction on $\kappa$, we can prove $\kappa\in W$. Since $W$ is closed under inhabited successor, and $\kappa\in W$, and $\kappa^{+}$ is inhabited, we have $\kappa^{+}\in W$. Since $\mathbb{F}$} is the intersection of all such sets $W$, and $\kappa^{+}$ belongs to every such $W$, we have $\kappa^{+}\in{\mathbb{F}}$ as desired. ∎
@@
$\emptyset\in{\sf zero}$   by the definition of $\sf zero$}
${\sf zero}\not\in{\sf zero}$   since $\sf zero$$\{\emptyset\}$ and $\sf zero$$\neq\emptyset$
$\emptyset\cup\{{\sf zero}\}\in{\sf zero}^{+}$   by definition of successor
@@
$\forall x\in{\sf FINITE}\,(|x|\in{\mathbb{F}}).$
Proof.
The formula to be proved is stratified, so we can prove it by induction on finite sets.
Base case: By Lemma 4.14, $|\emptyset|={\sf zero}$. By Lemma 4.19, ${\sf zero}\in{\mathbb{F}}$.
Induction step: Let $x\in{\sf FINITE}$ and $c\not\in x$. Consider $|x\cup\{c\}|$, which by Lemma 4.13 is $|x|^{+}$. By the induction hypothesis, $|x|\in{\mathbb{F}}$. By definition of $\mathbb{F}$}, $|x|^{+}\in{\mathbb{F}}$. That completes the induction step. ∎
@@
Proof.
By induction we prove
$\forall m\,(m\in{\mathbb{F}}\ {\rightarrow}\ \exists u\,(u\in m)).$
The formula is stratified, giving $u$ index 0 and $m$ index 1. For the base case, $\sf zero$$\{\emptyset\}$ by definition, so $\sf zero$} is inhabited. For the induction step, we always suppose $m^{+}$ is inhabited, so there is nothing more to prove. ∎
To put the proof directly: the set of inhabited members of $\mathbb{F}$} contains $\sf zero$} and is closed under inhabited successor, so it contains $\mathbb{F}$}.
@@
Proof.
Let $a$ be finite and $a\sim b$. Let $\kappa=|a|$. Then